# First principles band structure of interacting phosphorus and boron/aluminum $\delta$-doped layers in silicon

### v1

@@ Abstract
Silicon can be heavily doped with phosphorus in a single atomic layer (a $\delta$ layer), significantly altering the electronic structure of the conduction bands within the material. Recent progress has also made it possible to further dope silicon with acceptor-based $\delta$ layers using either boron or aluminum, making it feasible to create devices with interacting $\delta$ layers with opposite polarity. It is not known, however, how these $\delta$ layers will interact, particularly at small separation distances. Using Density Functional Theory, we calculate the electronic structure of a phosphorus-based $\delta$ layer interacting with a boron or aluminum $\delta$ layer, varying the distances between the $\delta$ layers. At separations 10 Å and smaller, the dopant potentials overlap and largely cancel each other out, leading to an electronic structure closely mimicking bulk silicon. At separations greater than 10 Å, the two layers behave independently of one another, forming a p-n diode with an intrinsic layer taking the place of the depletion region. One mechanism for charge transfer between $\delta$ layers at larger distances could be tunneling, where we see a greater than 3% probability for tunneling between a phosphorus and boron layer at 20 Å separation. This tunneling rate exceeds what would be seen for a standard silicon 1.1 eV triangular barrier, indicating that the interaction between delta layers creates enhanced tunneling at larger separation distances compared to a traditional junction. These calculations provide a foundation for the design of silicon-based electronics based on interacting $\delta$ layers.
@@ I Introduction
There are, however, relatively few explorations of the electronic structure of silicon subject to both $n$- and $p$-type APAM doping. Škereň et al. (2020); Mendez et al. (2025) For the creation of devices based on bipolar stacks of $\delta$ layers, it is important to understand how the electronic structure of these $\delta$ layers would interact, particularly at small separation distances. One intriguing possibility is that the induced conduction bands near the center of the Brillouin zone from the P $\delta$-layer will have their charge compensated by the induced valence bands from the B $\delta$-layer, moving the Fermi level into the band gap and creating a direct gap semiconductor within silicon. Conversely, the relevant defect potentials may interfere with one another, producing an electronic structure more closely resembling that of traditional silicon. This possibility strongly affects the understanding of tunneling in bipolar devices, where it is important to understand the conditions that enhance or suppress tunneling.
@@
In this manuscript, we use DFT to predict the electronic structure and local density of states of interacting phosphorus and boron/aluminum $\delta$-doped layers in silicon. We use the strongly constrained and appropriately normed (SCAN) exchange-correlation functionalSun et al. (2015) as a reasonable compromise between the computational cost of more accurate treatments of screened exchange (e.g., hybrid functionals) and the inaccuracy of more simplistic semilocal treatments. Our calculations predict that each layer creates a $\delta$ potential that hosts an impurity band, with the opposite charges compensating and mutually reducing the binding energies relative to the band edges. The extent to which the binding energies of the impurity bands are reduced is a function of the distance between the $\delta$ layers. At smaller separation distances << 10 Å, the two $\delta$-layers largely compensate each other and the electronic structure most closely resembles that of bulk silicon, making a direct gap unlikely. At distances >> 10 Å, the two layers become essentially independent. At these larger distances, the $\delta$ layers form a p-n junction with an intrinsic layer of silicon comprising the depletion region. We use the local potentials calculated within DFT to estimate the interlayer tunneling probability as a function of the separation between layers. It is enhanced relative to the tunneling probability for a triangular barrier with a height of the intrinsic silicon band gap (1.1 eV), indicating that interactions between the $\delta$ layers do lead to outcomes that would not be achieved with traditional diodes.
@@ II Methods
Figure 1: (a) The supercell used throughout this manuscript, indicating the $P$ and $B$ $\delta$ layers and the distance separating them. The horizontal axis corresponds to the (100) direction. (b) The band structure of a pure silicon supercell. Because we are using a supercell, the more familiar Brillouin zone for a two-atom primitive cell is folded in on itself. The red line represents the Fermi level of the system. (c) The Brillouin zone for the supercell, labeling high-symmetry points with crystal reciprocal coordinates. The kx axis, associated with the (001) direction of the supercell is exaggerated for clarity. Given the significant length of the axis in real space, this dimension is essentially negligible in reciprocal space.
All electronic structure calculations are done using the quantum espresso package.Giannozzi et al. (2009) We use the SCANSun et al. (2015) approximation to the exchange-correlation functional as implemented by Yao and Kanai.Yao and Kanai (2017) As shown in Fig. 1b, we predict a band gap of 0.98 eV for a pure silicon supercell. This is only slightly below the experimental valueBludau et al. (1974) of $\approx$1.1 eV, suggesting that estimates of other features of the band structure might achieve a similar degree of accuracy. We use kinetic energy cutoffs of 50 Ry and 200 Ry for the plane-wave basis sets used to describe the Kohn-Sham orbitals and charge density, respectively. We use a $\times2$$\times$1 Monkhorst-Pack gridMonkhorst and Pack (1976) to sample the Brillioun zone in our initial self-consistent calculation and then a $\times4$$\times$1 Monkhorst-Pack grid for non self-consistent calculations before band structures are calculated. We reference all energies to the conduction band minimum (CBM) of the host silicon for any given structure. Unless otherwise noted, the atomic positions are relaxed according to Born-Oppenheimer forces and considered converged below 50 meV/Å. The cell size is kept fixed, mimicking the embedding of the $\delta$-doped layer within the larger silicon structure. Band structures and local density of states (LDOS) are then calculated for each structure.
@@ III.1 Band structure and LDOS of boron-phosphorus $\delta$-doped layers
Figure 2: The band structure and local density of states (LDOS) (i.e. the localized band structure) for boron and phosphorus $\delta$-doped layers separated by (a) 1 Å, and (b) 4 Å. As the separation distance increases, the $\delta$-layer potentials become more visible within the LDOS, decreasing the band gap of the overall structure. The red line represents the Fermi level of the system.
@@
The band structure of a pair of phosphorus and boron $\delta$ layers separated by a single atomic layer ($1.24$ Å) is shown in Fig. 2a. At this range the layers compensate each other, leaving the material behaving largely as bulk silicon. The calculated band gap of the system is reduced slightly from 0.98 eV to 0.86 eV, although it should be noted that compressing the band diagram of two spatially separated $\delta$ layers can be somewhat misleading. We can see in the LDOS that the donor and acceptor $\delta$ potentials are spatially separated within the larger structure. Nonetheless, simulations of the electronic transport in bipolar devices have shown that band gap narrowing likely plays a significant role in the device behavior, Mendez et al. (2025) so we still report reduced band gap figures within this work. Given the strong charge screening at this short separation, the $\delta$ potentials more closely resemble two adjacent triangular potentials. This separation distance is likely too small to be precisely achieved experimentally, but might be approximated through a sequence of dosing a silicon surface with one dopant precursor, annealing to incorporate it, and repeating the sequence with a different precursor without depositing silicon between dosing sequences.
@@
At a separation of 4 Å, the $\delta$ potentials become much less suppressed, as seen in the band structure and LDOS in Fig. 2b. This increase in the prominence of the $\delta$-doped layers can be attributed to the intervening silicon layers between the donors and acceptors screening the charge from each other. The band gap of the total system is further reduced to 0.52 eV. Furthermore, the $\delta$ potentials are now clearly visible in the LDOS, showing a distinct protrusion for the phosphorus $\delta$ layer in the conduction band, and a lower magnitude, but still distinct protrusion in the valence band from the boron $\delta$ layer.
@@
Figure 3: The band structure and local density of states (LDOS) for boron and phosphorus $\delta$-doped layers separated by (a) 10 Å, (b) 20 Å, and (c) 100 Å. At these larger separation distances, the $\delta$-layer potentials become clearly distinct and, in the case of a 100 Å separation, overlapping in energy. The red line represents the Fermi level of the system.
@@
At separations of 10 and 20 Å, shown in Fig. 3a and b, respectively, the different $\delta$ layers become clearly distinct. At 10 Å, the induced VBM from the boron $\delta$-doped layer becomes clearly visible, shifting the Fermi level of the material within the valence bands. This is reflected in a band gap of 0.45 eV at 10 Å, and clearly visible $\delta$ potentials in the LDOS in both the conduction and valence band. This trend is continued at 20 Å, with a band gap of 0.36 eV and slightly less sharp $\delta$ potentials in the LDOS. The two $\delta$-doped layers are still interacting and suppressing the expression of each other, but this suppression reduces with separation, thus explaining the gradually decreasing band gap.
@@
At a distance of 100 Å, however, the two $\delta$-doped layers are fully independent. This leads to the induced valence and conduction bands from the $\delta$-doped layers overlapping in the middle of the band gap, as shown in Fig. 3. This is in line with what we would expect from an independent phosphorus $\delta$-doped layer with a magnitude $\approx$ 0.6 eVCarter et al. (2009) superimposed on an independent boron $\delta$-doped layer with a magnitude of $\approx$ 0.4 eV Campbell et al. (2023) away from the CBM and VBM, respectively. It should be cautioned, however, that while the overall band structure for this separation looks like a metal, examining the LDOS shows that the valence and conduction band peaks are spatially confined to the $\delta$-layer positioning and thus unlikely to directly overlap. Thus, this structure at 100 Å layer spacing is more akin to an ultra-shallow p-n junction, with a level of precision in dopant placement not attainable through traditional methods such as ion implantation.
@@
Figure 4: The band structure and local density of states (LDOS) for aluminum and phosphorus $\delta$-doped layers separated by (a) 1 Å, and (b) 4 Å. The aluminum atoms have less suppression of the $\delta$ doped layer induced bands, resulting in overlap even at these lower separation distances.
@@
Figure 5: The band structure and local density of states (LDOS) for aluminum and phosphorus $\delta$-doped layers separated by (a) 10 Å  (b) 20 Å and (c) 100 Å. The aluminum atoms have less suppression of the $\delta$ doped layer induced bands, resulting in overlap even at these lower separation distances.
@@ III.2 Band structure and LDOS of aluminum-phosphorus $\delta$-doped layers
In addition to the use of boron as an acceptor for $\delta$-doping a layer of silicon, aluminum has also been demonstrated as a potential dopant.Radue et al. (2021) We examine how these aluminum-phosphorus structures may compare with boron-phosphorus structures in Fig. 4 and 5, looking at the band structure and the LDOS of the material. We see that in these systems the $\delta$-doped layer induced bands are less suppressed by each other than the structures doped with phosphorus and boron, leading to more strongly overlapping band structures. In appendix B, we explore the doping potentials that are generated by these boron-phosphorus and aluminum-phosphorus structures. We demonstrate that boron induces more stress in the surrounding silicon than aluminum does. This observation can be used to help explain the significantly lowered amount of suppression of the $\delta$-layer induced bands with aluminum compared to boron. Stronger displacement from the boron $\delta$-doped layers can lead to small atomic variation in the position of nearby silicon atoms, which can help screen the potential from the $\delta$-doped layers from extending as far. Overall, however, these structures largely resemble the same band gap behavior seen in the boron and phosphorus $\delta$-doped layer structures, with effective band gaps lowering as the separation increases. We hypothesize that an arsenic-boron system or arsenic-aluminum system (or additional combinations moving further down the periodic table) would continue the trend of lowered suppression of the $\delta$-layer induced bands.
@@ III.3 Carrier transport through tunneling between $\delta$-doped layers
Given the spatially separated nature of the $\delta$-layer peaks, charge carrier movement between the layers (absent the application of any bias on the system) will likely be dominated by tunneling. The cost of an electron tunneling to the opposite polarity $\delta$-doped layer is dominated by overcoming the potential barrier induced by the interacting structure. We can calculate this tunneling probability of an electron sitting with energy $E$ at the phosphorus $\delta$-doped layers moving to the boron $\delta$-doped layer usingLandau and Lifshitz (2013)
@@
$k(z)=\sqrt{\dfrac{2m^{*}}{\hbar^{2}}(V(z)-E)}.$ (2)
Here, $m^{*}$ is the effective mass of the electron, and $V(z)$ is the potential of the system as a function of the $z$, position. We can conveniently abstract a $V(z)$ function from our DFT LDOS calculations, and numerically integrate to calculate the tunneling probability in our systems. We then compare this to the tunneling that would be expected for a pure triangular 1.1 eV barrier in Fig. 6.
Notably, due to the interaction between the two $\delta$-doped layers, the tunneling probability remains higher than would be expected for a pure triangular barrier. This trend is particularly pronounced for longer separation distances. We predict the tunneling probability of a boron and phosphorus $\delta$-doped layer separated by 10 and 20 Å is 0.141 and 0.0305, respectively. While these are not large probabilities, they are significant enough that a reasonable number of electrons may tunnel through. The tunneling rate in aluminum and phosphorus $\delta$-doped structures are lower than for boron and phosphorus structures due to the stronger peaks induced by less relaxation of nearby silicon. This trend continues throughout the system, however, the slope of the Al-P $\delta$ layer tunneling remains above that of a 1.1 eV triangular barrier. We note that these results are before any potential bias is applied to the system, which could be used to further manipulate tunneling rates. This tunneling probability calculation is based on an electron sitting at the bottom of the phosphorus $\delta$-layer potential well and assuming that it can tunnel to the top of the acceptor $\delta$ well. This assumption of overlapping energy levels in the donor and acceptor is clearly true at high separation distances, but seems inaccurate for separations below 10 Å. This is particularly true for boron $\delta$ layers, where the Fermi level does not sit between the valence and conduction bands, but seems to be caught within the B $\delta$ layer. This means that the easiest electron movement within the system would still be within the $\delta$ layer, making the behavior of the material metallic. We next examine how easy it is to move electrons around the Fermi level of the system.
@@
The density of states (DOS) at the Fermi level of the system gradually increases as the separation between the boron and phosphorus $\delta$-layers increases, as shown in Fig. 7. At lower separations, the DOS is quite low, as the $\delta$-layer induced bands are suppressed and the Fermi level is located within the band gap of the larger structure. As the separation between the delta layers increases, however, the $\delta$-layer induced valence and conduction bands increase in size and the Fermi level moves from the band gap into the valence bands. This causes the DOS at the Fermi level to increase with the separation between the materials until they appear to saturate at 10 Å separation. These higher separations are thus more likely to induce weak metallic-like behavior as electrons can easily move between bands in the valence band structure. While the aluminum-phosphorus structures have systematically lower density of states, the trends remain the same.

### v2

@@ Abstract
Silicon can be heavily doped with phosphorus in a single atomic layer (a $\delta$ layer), significantly altering the electronic structure of the conduction bands within the material. Recent progress has also made it possible to further dope silicon with acceptor-based $\delta$ layers using either boron or aluminum, making it feasible to create devices with interacting $\delta$ layers with opposite polarity. Using Density Functional Theory, we calculate the electronic structure of a phosphorus-based $\delta$ layer interacting with a boron or aluminum $\delta$ layer, varying the distances between the $\delta$ layers. At separations 1 nm and smaller, the dopant potentials overlap and largely cancel each other out, leading to an electronic structure closely mimicking intrinsic silicon. At separations greater than 1 nm, the two $\delta$ layers behave independently of one another, with an equivalent electronic structure to a p-n diode with an intrinsic layer taking the place of the depletion region. One mechanism for charge transfer between $\delta$ layers at larger distances could be tunneling, where we see a tunneling probability exceeding what would be seen for a standard silicon 1.1 eV triangular barrier, indicating that the interaction between delta layers may enhance tunneling compared to a traditional junction.
preprint: AIP/123-QED
@@ I Introduction
There are, however, relatively few explorations of the electronic structure of silicon subject to both $n$- and $p$-type APAM doping. Škereň et al. (2020); Mendez et al. (2025) For the creation of devices based on bipolar stacks of $\delta$ layers, it is important to understand how the electronic structure of these $\delta$ layers would interact, particularly at small separation distances. One intriguing possibility is that the induced conduction bands near the center of the Brillouin zone from the P $\delta$-layer will have their charge compensated by the induced valence bands from the B $\delta$-layer, moving the Fermi level into the band gap and creating a direct gap semiconductor within silicon. Conversely, the relevant defect potentials may interfere with one another, producing an electronic structure more closely resembling that of intrinsic silicon. This possibility strongly affects the understanding of tunneling in bipolar devices, where it is important to understand the conditions that enhance or suppress tunneling.
@@
In this manuscript, we use DFT to predict the electronic structure and local density of states of interacting phosphorus and boron/aluminum $\delta$-doped layers in silicon, building upon previous work that has only looked at a single type of $\delta$-doped layer. We use the strongly constrained and appropriately normed (SCAN) exchange-correlation functionalSun et al. (2015) as a reasonable compromise between the computational cost of more accurate treatments of screened exchange (e.g., hybrid functionals) and the inaccuracy of more simplistic semilocal treatments, again providing novel results. Our calculations predict that each layer creates a $\delta$ potential that hosts an impurity band, with the opposite charges compensating and mutually reducing the binding energies relative to the band edges. The extent to which the binding energies of the impurity bands are reduced is a function of the distance between the $\delta$ layers. At smaller separation distances << 1 nm, the two $\delta$-layers largely compensate each other and the electronic structure most closely resembles that of intrinsic silicon, making a direct gap unlikely. At distances >> 1 nm, the two layers become essentially independent. At these larger distances, the $\delta$ layers essentially form the electronic structure of a p-n junction with an intrinsic layer of silicon comprising the depletion region. We use the local potentials calculated within DFT to estimate the interlayer tunneling probability as a function of the separation between layers. It is enhanced relative to the tunneling probability for a triangular barrier with a height of the intrinsic silicon band gap (1.1 eV), indicating that interactions between the $\delta$ layers do lead to outcomes that would not be achieved with traditional diodes.
@@ II Methods
Figure 1: (a) The supercell used throughout this manuscript, indicating the $P$ and $B$ $\delta$ layers and the distance separating them. The horizontal axis corresponds to the (100) direction. (b) The band structure of a pure silicon supercell. Because we are using a supercell, the more familiar Brillouin zone for a two-atom primitive cell is folded in on itself. The red line represents the Fermi level of the system. (c) The Brillouin zone for the supercell, labeling high-symmetry points with crystal reciprocal coordinates. The length of the box in the kx direction, associated with the (001) direction of the supercell is exaggerated for clarity. Given the significant length of the axis in real space, this dimension is essentially negligible in reciprocal space.
All electronic structure calculations are done using the quantum espresso package.Giannozzi et al. (2009) We use the SCANSun et al. (2015) approximation to the exchange-correlation functional as implemented by Yao and Kanai.Yao and Kanai (2017) As shown in Fig. 1b, we predict a band gap of 0.98 eV for a pure silicon supercell. This is only slightly below the experimental valueBludau et al. (1974) of $\approx$1.1 eV, suggesting that estimates of other features of the band structure might achieve a similar degree of accuracy. We use kinetic energy cutoffs of 680 eV and 2721 eV for the plane-wave basis sets used to describe the Kohn-Sham orbitals and charge density, respectively. We use a $\times2$$\times$1 Monkhorst-Pack gridMonkhorst and Pack (1976) to sample the Brillioun zone in our initial self-consistent calculation and then a $\times4$$\times$1 Monkhorst-Pack grid for non self-consistent calculations before band structures are calculated. We reference all energies to the conduction band minimum (CBM) of the intrinsic silicon for any given structure. Unless otherwise noted, the atomic positions are relaxed according to Born-Oppenheimer forces and considered converged below 0.5 eV/nm. The cell size is kept fixed, mimicking the embedding of the $\delta$-doped layer within the larger silicon structure. Band structures and local density of states (LDOS) are then calculated for each structure.
@@ III.1 Band structure and LDOS of boron-phosphorus $\delta$-doped layers
Figure 2: The band structure and local density of states (LDOS) (i.e. the localized band structure) for boron and phosphorus $\delta$-doped layers separated by (a) 0.1 nm, and (b) 0.4 nm. As the separation distance increases, the $\delta$-layer potentials become more visible within the LDOS, decreasing the band gap of the overall structure. The red line represents the Fermi level of the system.
@@
The band structure of a pair of phosphorus and boron $\delta$ layers separated by a single atomic layer ($0.124$ nm) is shown in Fig. 2a. At this range the layers compensate each other, leaving the material behaving largely as intrinsic silicon. The calculated distance from the top of the valence band to next band higher in energy is reduced slightly from 0.98 eV to 0.86 eV. It should be noted that the band diagram of two spatially separated $\delta$ layers can be somewhat misleading, as the z direction of the Brillouin zone is highly compressed. While the band structure appears to show a direct gap, we can see in the LDOS that the donor and acceptor $\delta$ potentials are spatially separated within the larger structure. Nonetheless, simulations of the electronic transport in similarly spatially separated bipolar devices have shown that band gap narrowing likely plays a significant role in the resulting device behavior, Mendez et al. (2025) so we still report on the energetic separation between the donor and acceptor $\delta$ bands in figures within this work. We avoid directly referring to this as a band gap since an electron would typically still need $\approx$1 eV energy to move directly from the valence band to the conduction band, according to the LDOS shown in Fig. 2a. Given the strong charge screening at this short separation, the $\delta$ potentials more closely resemble two adjacent triangular potentials in the corresponding LDOS diagrams. This separation distance is likely too small to be precisely achieved experimentally, but might be approximated through a sequence of dosing a silicon surface with one dopant precursor, annealing to incorporate it, and repeating the sequence with a different precursor without depositing silicon between dosing sequences.
@@
At a separation of 0.4 nm, the $\delta$ potentials become much less suppressed, as seen in the band structure and LDOS in Fig. 2b. This increase in the prominence of the $\delta$-doped layers can be attributed to the intervening silicon layers between the donors and acceptors screening the charge from each other. The energetic distance from the boron $\delta$ layer peak to the phosphorus $\delta$ layer peak further narrows to 0.52 eV. Furthermore, the $\delta$ potentials are now clearly visible in the LDOS in Fig. 2b, showing a distinct protrusion for the phosphorus $\delta$ layer in the conduction band, and a lower magnitude, but still distinct protrusion in the valence band from the boron $\delta$ layer.
@@
Figure 3: The band structure and local density of states (LDOS) for boron and phosphorus $\delta$-doped layers separated by (a) 1 nm, (b) 2 nm, and (c) 10 nm. At these larger separation distances, the $\delta$-layer potentials become clearly distinct and, in the case of a 10 nm separation, overlapping in energy. The red line represents the Fermi level of the system.
@@
At separations of 1 and 2 nm, shown in Fig. 3a and b, respectively, the different $\delta$ layers become clearly distinct in the resulting LDOS diagrams. At 1 nm, the induced VBM from the boron $\delta$-doped layer becomes clearly visible, shifting the Fermi level of the material within the valence bands. This is reflected in the energy difference from the boron $\delta$ layer peak to the phosphorus $\delta$ layer peak shifting to 0.45 eV at 1 nm, and clearly visible $\delta$ potentials in the LDOS in both the conduction and valence band. This trend is continued at 2 nm, with an energy difference between the boron $\delta$ layer peak to the phosphorus $\delta$ layer peak of 0.36 eV and slightly less sharp $\delta$ potentials in the LDOS. The two $\delta$-doped layers are still interacting and suppressing the expression of each other, but this suppression reduces with separation, thus explaining the gradually decreasing energy between the boron and phosphorus $\delta$ layer peaks.
@@
At a distance of 10 nm, however, the two $\delta$-doped layers are fully independent. This leads to the induced valence and conduction bands from the $\delta$-doped layers overlapping in the middle of the band gap, as shown in Fig. 3. This is in line with what we would expect from an independent phosphorus $\delta$-doped layer with a magnitude $\approx$ 0.6 eVCarter et al. (2009) superimposed on an independent boron $\delta$-doped layer with a magnitude of $\approx$ 0.4 eV Campbell et al. (2023) away from the CBM and VBM, respectively. It should be cautioned, however, that while the overall band structure for this separation looks like a metal, examining the LDOS shows that the valence and conduction band peaks are spatially confined to the $\delta$-layer positioning and thus unlikely to directly overlap. Thus, this structure at 10 nm layer spacing is more akin to an ultra-short p-n junction, with a level of precision in dopant placement not attainable through traditional methods such as ion implantation.
@@
Figure 4: The band structure and local density of states (LDOS) for aluminum and phosphorus $\delta$-doped layers separated by (a) 0.1 nm, and (b) 0.4 nm. The aluminum atoms have less suppression of the $\delta$ doped layer induced bands, resulting in overlap even at these lower separation distances.
@@
Figure 5: The band structure and local density of states (LDOS) for aluminum and phosphorus $\delta$-doped layers separated by (a) 1 nm, (b) 2 nm and (c) 10 nm. The aluminum atoms have less suppression of the $\delta$ doped layer induced bands, resulting in overlap even at these lower separation distances.
@@ III.2 Band structure and LDOS of aluminum-phosphorus $\delta$-doped layers
In addition to the use of boron as an acceptor for $\delta$-doping a layer of silicon, aluminum has also been demonstrated as a potential dopant.Radue et al. (2021) We examine how these aluminum-phosphorus structures may compare with boron-phosphorus structures in Fig. 4 and 5, looking at the band structure and the LDOS of the material. We see that in these systems the $\delta$-doped layer induced bands in the band structure are less suppressed by each other than the structures doped with phosphorus and boron, leading to more strongly overlapping band structures. In appendix B, we explore the doping potentials that are generated by these boron-phosphorus and aluminum-phosphorus structures. We demonstrate that boron induces more stress in the surrounding silicon than aluminum does. This observation can be used to help explain the significantly lowered amount of suppression of the $\delta$-layer induced bands with aluminum compared to boron, and generally cleaner valence and conduction peaks in the resulting LDOS. Stronger displacement from the boron $\delta$-doped layers can lead to small atomic variation in the position of nearby silicon atoms, which can help screen the potential from the $\delta$-doped layers from extending as far. Overall, however, these structures largely resemble the same behavior seen in the boron and phosphorus $\delta$-doped layer structures, with the energy between the acceptor and donor peaks lowering as the separation increases. The LDOS for these systems also remain similar, with distinct acceptor and donor peaks forming at $\delta$ layer separation distances >> 1 nm. We hypothesize that an arsenic-boron system or arsenic-aluminum system (or additional combinations moving further down the periodic table) would continue the trend of lowered suppression of the $\delta$-layer induced bands.
@@ III.3 Carrier transport through tunneling between $\delta$-doped layers
Given the spatially separated nature of the $\delta$-layer peaks, if there is any charge carrier movement between the layers (absent the application of any bias on the system), one avenue is tunneling. The cost of an electron tunneling to the opposite polarity $\delta$-doped layer is dominated by overcoming the potential barrier induced by the interacting structure. We can approximate the tunneling probability of a single electron sitting with energy $E$ at the phosphorus $\delta$-doped layers moving to the boron $\delta$-doped layer usingLandau and Lifshitz (2013)
@@
[…]
Here, $m^{*}$ is the effective mass of the electron, and $V(z)$ is the potential of the system as a function of the $z$, position. We assume an effective mass of $m^{*}=0.19m_{e}$, where $m_{e}$ is the standard mass of an electron, matching the standard transverse mass of an electron in silicon, and the tunneling is solved for a single electron wavefunction potential landscape $V(z)$. We can conveniently abstract a $V(z)$ function from our DFT LDOS calculations (which are shown in Appendix C), and numerically integrate to calculate the tunneling probability in our systems. We then compare this to the tunneling that would be expected for a pure triangular 1.1 eV barrier in Fig. 6.
Notably, due to the interaction between the two $\delta$-doped layers, the tunneling probability remains higher than would be expected for a pure triangular barrier. This trend is particularly pronounced for longer separation distances. We predict the tunneling probability of a boron and phosphorus $\delta$-doped layer separated by 1 and 2 nm is 0.141 and 0.0305, respectively. While these are not large probabilities, they are significant enough that a reasonable number of electrons may tunnel through. The tunneling rate in aluminum and phosphorus $\delta$-doped structures are lower than for boron and phosphorus structures due to the stronger peaks induced by less relaxation of nearby silicon. This trend continues throughout the system, however, the slope of the Al-P $\delta$ layer tunneling remains above that of a 1.1 eV triangular barrier. We note that these results are before any potential bias is applied to the system, which could be used to further manipulate tunneling rates. This tunneling probability calculation is based on an electron sitting at the bottom of the phosphorus $\delta$-layer potential well and assuming that it can tunnel to the top of the acceptor $\delta$ well. This assumption of overlapping energy levels in the donor and acceptor is clearly true at high separation distances, but seems inaccurate for separations below 1 nm. This is particularly true for boron $\delta$ layers, where the Fermi level does not sit between the valence and conduction bands, but seems to be caught within the B $\delta$ layer. We therefore show these tunneling results merely to illustrate the possibility of an electron moving between layers, and compare how easy it would be in this system compared to a generic triangular potential well. The easiest electron movement within the given system would likely still be within a $\delta$ layer, making the behavior of the material essentially metallic around the delta layer. We next examine how easy it is to move electrons around the Fermi level of the system.
@@
The density of states (DOS) at the Fermi level of the system gradually increases as the separation between the boron and phosphorus $\delta$-layers increases, as shown in Fig. 7. At lower separations, the DOS is quite low, as the $\delta$-layer induced bands are suppressed and the Fermi level is located within the band gap of the larger structure. As the separation between the delta layers increases, however, the $\delta$-layer induced valence and conduction bands increase in size and the Fermi level moves from the band gap into the valence bands. This causes the DOS at the Fermi level to increase with the separation between the materials until they appear to saturate at 1 nm separation. These higher separations are thus more likely to induce weak metallic-like behavior as electrons can easily move between bands in the valence band structure. While the aluminum-phosphorus structures have systematically lower density of states, the trends remain the same.